OpenStudy (lucaz):

f(x)=x(x-4)(x+4) and g(x)=(x²-36)(x²-4) how do I find the values for f(x) x g(x) > 0 ?

OpenStudy (amistre64):

define where f or g equals zero, then run some test points inbetween the intervlas

OpenStudy (amistre64):

the intervals that are + are by default ... greater than zero

OpenStudy (lucaz):

I have the zeros of both

OpenStudy (amistre64):

then lets test them out ....

OpenStudy (amistre64):

fg = x(x-4)(x+4)(x²-36)(x²-4) -6 -4 -2 0 2 4 6 let x = +- 1, +- 3, +- 5 and +- 10 and see how they interact

OpenStudy (amistre64):

for x=-1 x (x-4) (x+4) (x²-36) (x²-4) - - + - - -- = + -- = + + = + therefore the overall sign is + between -2 and 0

OpenStudy (lucaz):

oh, I see..

OpenStudy (lucaz):

thank you

OpenStudy (amistre64):

youre welcome